Explain energy of a capacitor?
1
Expert's answer
2020-10-05T10:58:14-0400

The capacitor energy is the potential energy of the interaction of the capacitor plates - after all, the plates, being charged differently, are attracted to each other.

Take such a small platform on the second plate of the capacitor that the charge "q_0" this site can be considered point. This charge is attracted to the first lining with force "F_0 = q_0E_1" , where "E_1" is the field strength of the first lining:

"E_1=\frac{\displaystyle \sigma }{\displaystyle 2 \varepsilon _0 \vphantom{1^a}}=\frac{\displaystyle q}{\displaystyle 2\varepsilon_0 S \vphantom{1^a}}."

Therefore,

"F_0=\frac{\displaystyle q_0q}{\displaystyle 2 \varepsilon_0 S \vphantom{1^a}}."

This force is directed parallel to the field lines (i.e. perpendicular to the plates).

The resulting attraction force "F" of the second cover to the first cover is composed of all these forces "F_0" with which all kinds of small charges "q_0" the second cover are attracted to the first cover. In this summation, the constant multiplier "\frac q {2\varepsilon _ 0 S}" is placed outside the bracket, and in the bracket all "q_0" are summed and "q" is given. As a result, we get:

"F=\frac{\displaystyle q^2}{\displaystyle 2 \varepsilon_0 S \vphantom{1^a}}" . (1)

Suppose now that the distance between the plates has changed from the initial value "d_1" to the final value "d_2" . The force of attraction of the plates performs the following work:

"A = F(d_1 - d_2)" .

The sign is correct: if the plates approach "(d_2 < d_1)" , then the force performs positive work, since the plates are attracted to each other. On the contrary, if you remove the plates "(d_2 > d_1)" , then the operation of the attraction force is negative, as it should be.

Taking into account the formulas (1) and "C=\frac{\displaystyle \varepsilon_0 S}{\displaystyle d \vphantom{1^a}}" - formula of flat air condenser capacity, we have:


"A=\frac{\displaystyle q^2}{\displaystyle 2 \varepsilon_0 S \vphantom{1^a}}\left ( d_1-d_2 \right )=\frac{\displaystyle q^2d_1}{\displaystyle 2\varepsilon_0 S \vphantom{1^a}}-\frac{\displaystyle q^2d_2}{\displaystyle 2\varepsilon_0 S \vphantom{1^a}}" "=\frac{\displaystyle q^2}{\displaystyle 2C_1 \vphantom{1^a}}-\frac{\displaystyle q^2}{\displaystyle 2C_2 \vphantom{1^a}}=W_1-W_2,"

where

"W_1=\frac{\displaystyle q^2}{\displaystyle 2C_1 \vphantom{1^a}},"

"W_2=\frac{\displaystyle q^2}{\displaystyle 2C_2 \vphantom{1^a}}"

This can be rewritten as follows:

"A = -(W_2 - W_1) = - \Delta W,"

where

"W=\frac{\displaystyle q^2}{\displaystyle 2C \vphantom{1^a}}" .

The operation of the potential attraction force "F" of the plates turned out to be equal to a change with a sign minus the value of "W". This just means that "W" is the potential interaction energy of the plates, or the energy of a charged capacitor.